Shift summary - Owl 11/10/99

Arun Tripathi (arun@physics.ucla.edu)
Wed, 10 Nov 1999 05:01:25 -0800 (PST)


Shift Summary for Owl shift, 10/Nov/99

On Shift: Carlos Escobar, Arun Tripathi

Hours of beam available at all intensities: 7
Hours of beam available at nominal intensity: 7
Hours of E799 data (special runs not included): 6 hrs 15min
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Initializing pipeline to fix the laser
illumination problem, stopping the run
due to full tapes, changing tapes, kfrending.

========================================================================
Runs ended this shift:
Run 14937 - 10.9M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

-------------------------------------------------------------------------
Accesses during shift: None.

========================================================================
Overview of shift:
Relatively smooth data taking. After the tapes filled up, the run
won't stop gracefully, had to do a fresh start.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14937 93 CsI channel 2297 drops out of laser
illumination plot. Did hard boot of
pipeline cr11. Lost about 10 spills.

4:55 Tapes fill up. The run won't stop.
saved the l3 histograms, and did
drt_fresh_start. Kumquat 3x DYC
busy comes up. Required several kfrends
and reloading of the fruit.

14938 In progress, without any problems, so
far.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift